OpenStudy (anonymous):

a,b,c are complex numbers: prove: if |a|=|b|=|c|=1 then (conjugate of)(a+b+c)= 1/a+1/b+1/c

OpenStudy (jamesj):

If |a|=|b|=|c|=1, then a,b,c all lie on the unit circle an may be written as a = e^(pi) b = e^(qi) c = e^(ri) for real numbers, p, q, r. Then a* (conjugate) = e^(-pi) = (e^(pi))^-1 = 1/a, b* = 1/b and c* = 1/c. Therefore (a+b+c)* = ...